8. If a wave has a frequency of 440 Hz, how long does it take for 10° of a cycle to occur? (a) 0.00273 s (b) 0.000273 s (c) 0Can you show me how to do this step by step? Please write legible. Thank you!
Answer #1

wt = O = 10 = 10 xal) wt=0, 0-100 - 10 x 25 - I rad. 360 - t - 1 x en 18x21 18x2x4 40 = 0.0000631 t=0.0000 6315
Answer #2

SOLUTION :


Wave frequency = 440 Hz = 440 cycles per second


So, time needed for 1 cycle (360º)  = 1 / 440 sec.


So, time needed for 10º =  1 /440 * 10 / 360 = 0.0000631 sec = Option c (ANSWER).

answered by: Tulsiram Garg